OpenStudy (anonymous):

f(x)=2x^2−146x+c has 2 roots that are positive prime numbers. What is c?

OpenStudy (anonymous):

or roots to be positive D>0 i.e. b^2 - 4ac >0 D=b^2 - 4ac = (-146)^2 - (4*2*c)>0 =>21316 - 8c>0 =>8c<21316 => c<5329/2 is dis the ans??

Parth (parthkohli):

use the sum of roots and product of roots formulae.

Parth (parthkohli):

\[-\dfrac{146}{2} = \alpha + \beta\]and\[\dfrac{c}{a} = \alpha \beta\]where \(\alpha\), \(\beta\) are the required prime roots.

Parth (parthkohli):

Know any two prime numbers with sum \(73\)?

Parth (parthkohli):

I meant\[\dfrac{-(-146)}{2} = \alpha + \beta\]

OpenStudy (anonymous):

ya its 73

OpenStudy (anonymous):

but 2 prime number tat makes 73??

Parth (parthkohli):

Yeah, do you know any two such prime numbers with sum 73?

OpenStudy (klimenkov):

If one of this primes is odd, the second must be even, because \(\alpha +\beta=73\). But how many even primes do you know?

OpenStudy (anonymous):

Start from the very first prime number and work your way through. That's a hint for you.

Parth (parthkohli):

There is only one even prime number.

OpenStudy (anonymous):

I stress on the word "first" in that statement.

Parth (parthkohli):

And of course, when you subtract the even prime, you will get another prime.

OpenStudy (klimenkov):

@ParthKohli 17 - prime. 2 - even prime. If I subtract 17 - 2 = 15 - I will get not prime.

Parth (parthkohli):

No, in this case you must!

OpenStudy (anonymous):

This case is NOT having such a situation though :-)

Parth (parthkohli):

...?

Parth (parthkohli):

This question would have no answer if you would not get another prime number after subtracting 2.

OpenStudy (klimenkov):

Nobody guarantees you that this must have an answer.

OpenStudy (anonymous):

Hey this is a problem of the present week's Brilliant Challenges... Shouldn't be discussed here actually... Between, Awesome site for maths lovers... https://brilliant.org/

OpenStudy (shubhamsrg):

really? :O

OpenStudy (anonymous):

Yeah...

OpenStudy (shubhamsrg):

woah! :O

Parth (parthkohli):

LOL!!! ^^